OpenStudy (anonymous):

Use the method of substitution to solve the system. x^2+ y^2 = 8 y − x = 4

jimthompson5910 (jim_thompson5910):

y − x = 4 y = 4 + x

jimthompson5910 (jim_thompson5910):

x^2+ y^2 = 8 x^2+ (4+x)^2 = 8 x^2 + 16 + 8x + x^2= 8 2x^2 + 8x + 16 = 8 2x^2 + 8x + 16 - 8 = 0 2x^2 + 8x + 8 = 0 2(x^2 + 4x + 4) = 0 x^2 + 4x + 4 = 0 I'll let you finish up